# Assessing (im)balance in signed brain networks

### v1

@@ I.2 From univariate time series to graphs
Different (classes of) approaches exist that seek to extract a network from one or more time series while ignoring (the intermediate step represented by) the proximity matrix [^9] We redirect the interested reader to the reviews Silva et al. (2021); Zou et al. (2019), discussing the differences between the existing (classes of) algorithms and comparing their performances..
@@ I.2.1 Visibility-induced mapping
[…]
@@
A variant of the aforementioned procedure has been proposed in Luque et al. (2009), where the concept of horizontal visibility - as opposed to the one of natural visibility - is defined. A horizontal visibility graph (HVG) is constructed by considering (only) horizontal visibility lines: as a consequence, the HVG is always a subgraph of the VG, hence containing less information, although its construction is faster. HVGs have been employed to analyse seismic signals Telesca and Lovallo (2012), the volatility of returns of financial time series Zhang et al. (2015), heartbeat rates Xie et al. (2019) and for classification tasks Li et al. (2018).
@@
A third variant is defined by implementing the concept of directed visibility, whose introduction has been justified by noticing that time has a natural direction: in Lacasa et al. (2012), the authors adopted the rule according to which any pair of time steps $t_{i}<t_{j}$ populates a directed horizontal visibility graph (DHVG) with an edge $i\rightarrow j$ and employed DHVGs to quantify time series irreversibility.
@@
A fourth variant is defined by implementing the concept of limited penetrable visibility, proposed in Zhou et al. (2012) to reduce the level of noise accompanying data to a larger extent than that achievable by using VGs and HVGs. In this case, any two nodes are connected if the visibility line joining the tops of the two associated bars crosses a number of ‘obstacles’ that is smaller than, or equal to, a fixed value $L$ naturally, the limited penetrable visibility graph (LPVG) reduces to the VG when $L=0$, i.e. when there are no ‘obstacles’. LPVGs have been employed to analyse EEG signals of people in different conditions (affected by Alzheimer’s disease Wang et al. (2016) or under manual acupuncture Pei et al. (2014)).
@@
A fifth variant is defined by implementing the concept of weighted visibility. The method, proposed in Supriya et al. (2016), prescribes to enrich each edge of a DHVG with a weight matching the ‘view angle’ of the observing node, its sign with respect to the horizontal axis indicating if the values of the time series increase (if positive) or decrease (if negative); since such a method allows for negative weights, the absolute value of the angle is often considered. The authors employed the method to identify EEG signals associated with epilepsy.
@@
A last variant is defined by implementing the concept of differential visibility, realised by removing the edge set of the HVG from the edge set of the VG Zhu et al. (2014).
@@ I.2.2 Transition-induced mapping
The mapping introduced in Campanharo et al. (2011) is based on a partition of the distribution of the values assumed by a time series in quantiles, to each of which a symbol is assigned - the number of quantiles is an arbitrary parameter whose choice should depend on the length of the series itself Shirazi et al. (2009). The variant of this method introduced in Weng et al. (2017) allows ‘memory networks’ to be constructed by assigning the attribute $t$ to the edge directed from $i$ to $j$ if the transition between the corresponding symbols occurs at time step $t$. A second variant is the one introduced in Campanharo and Ramos (2016), where the concept of renewal intervals, i.e. transitions between symbols posed at various temporal distances, is introduced.
The mapping introduced in Gao and Jin (2009) is named Coarse-Grained Phase Space Graphs and requires the definition of a time window $\omega$ to determine the number of values of the series to be represented as an $\omega$-dimensional vector, whose components are numerically determined by adopting a rule similar to the quantiles-based one.
A different, yet related, mapping is the one introduced in Small (2013), named Ordinal Partition Transition Networks. Here, the time series is embedded in a $\omega$-dimensional space where, starting from each point of the series, the current and the following $\omega-1$ numbers, interspaced by a time delay $\tau$, are considered (for instance, setting $\omega=3$ and $\tau=2$, the $i$-th embedding vector will contain the values at $i$, $i+2$, and $i+4$). The embedding vectors are, then, associated with patterns induced by the rank of their values and each ordinal pattern is represented as a node of the network: nodes $i$ and $j$ are connected if the corresponding patterns occur at subsequent time steps (i.e. if $t_{j}=t_{i}+1$).
A third variant of the same kind of mapping is the one introduced in Stephen et al. (2015) and named Visibility Graphlets Networks. This method requires the definition of a time window $\omega$ to determine the number of values to be considered: the time steps from $t$ to $t+\omega$ are, then, mapped into a directed visibility graph (DVG). Such a procedure is repeated for each time step and the DVGs having the same adjacency matrix are associated with the same symbol, i.e. one of the nodes of the resulting network. Different nodes are linked according to the subsequent occurrences of the related DVGs.
@@ I.2.3 Proximity-induced mapping
The first method of this kind was introduced in Zhang et al. (2006) to represent pseudo-periodic time series. First, it segments a time series into cycles (e.g. dictated by a periodic behaviour) but without requiring them to have the same length; then, correlations between pairs of cycles are computed - by shifting the shortest one, if needed - and the largest correlation value is compared against a threshold: if found to be larger, an edge between the two cycles, represented as nodes in the resulting network, is established.
An alternative way of employing correlations to implement a proximity-induced mapping is introduced in Yang and Yang (2008) and prescribes to compute the correlations between the (either overlapping or non-overlapping) portions into which a time series can be partitioned: the nodes representing them are linked if the correlation between the corresponding portions exceeds a certain threshold.
A different class of methods is the one gathering the so-called Recurrence Networks (RNs) and is based on the idea that the points of a time series are recurrent if their distance, in a properly defined embedding space, is ‘short enough’. Similarly to the prescription characterising the Ordinal Partition Transition Networks algorithm, each time step is mapped into a point of an embedding space. Distances are, then, computed by using one of the many norms that can be defined. The way the embedding space is analysed distinguishes RNs in three sub-categories: i) the first one, introduced in Small et al. (2009), is composed by the so-called k-Nearest Neighbour Networks and requires to connect each node with the $k$ nodes whose corresponding points in the embedding space are closest to its corresponding point - to be noticed that such a variant has been defined for directed networks; ii) the second one, introduced in Xu et al. (2008), is composed by the so-called Adaptive Nearest Neighbour Networks and implements the undirected version of the aforementioned recipe; iii) the third one, introduced in Donner et al. (2010), is composed by the so-called $\epsilon$-Recurrence Networks ($\epsilon$-RNs) and prescribes to link the nodes whose corresponding points lie at a distance smaller than $\epsilon$: in symbols, the generic element of an $\epsilon$-RN matrix is defined as
@@ I.3.1 Single-layer mapping
This first kind of mapping collects generalisations of the aforementioned RNs, hereby named Cross Recurrence Networks (CRNs), Inter-System Recurrence Networks (ISRNs) and Joint Recurrence Networks (JRNs) Feldhoff et al. (2012, 2013); Romano et al. (2004).
@@
with $\bm{x}_{t}$ and $\bm{y}_{t^{\prime}}$ indicating the points of the chosen embedding space representing the values of the time series $X_{t}$ and $Y_{t^{\prime}}$ notice that the two time series do not need to be of the same length, hence inducing a rectangular table Marwan et al. (2002); Zbilut et al. (1998).
@@
constituted by (matrices representing the) RNs along the diagonal and by (matrices representing the) CRNs off the diagonal Feldhoff et al. (2012).
An alternative recipe is represented by the JRNs Romano et al. (2004); Feldhoff et al. (2013), inducing adjacency matrices like the one reading
@@
A generalisation of the Ordinal Partition Transition Networks for multivariate time series was introduced in Zhang et al. (2017). The method requires a binarisation of the univariate time series, letting an increasing trend [^10] Defined as ‘the value of the series at time $t$ is larger than the value of the series at time $t-1$’. being represented with a 111 and a decreasing trend being represented with a 0: upon doing so, $2^{N}$ combinations, with $N$ being the number of time series, are possible for each time step, $t$; finally, each combination is assigned a node in the resulting network, any two nodes connected by a directed edge representing combinations found at subsequent steps. For deterministic and periodic time series, some of the nodes remain isolated; random processes are, instead, associated with densely (if not completely) connected graphs.
@@ I.3.2 Multiple-layer mapping
This second kind of mapping, instead, associates a multivariate time series with a multi-layer network. A first example of such a mapping is provided by Lacasa et al. (2015): here, a multiplex visibility graph (MVG) is obtained by mapping each univariate time series into a graph by employing a VG; afterwards, the MVG is aggregated into a single-layer weighted graph, each node representing a layer-specific VG and each weight proxying the similarity between the corresponding layers. Such a method has been employed to analyse fMRI data Sannino et al. (2017).
The authors of Eroglu et al. (2018) follow the same pipeline but replace VGs with RNs, thus employing Multiplex Recurrence Networks (MRNs) to study paleoclimatic time series.
@@ I.4 From multivariate time series to
The most recent approaches emphasise the importance of probabilistic (null, as well as generative) models to benchmark network features and disentangle fundamental, structural properties from statistical artifacts Varoquaux and Craddock (2013); Váša and Mišić (2022). These models play a crucial role in understanding whether empirical features (e.g. a modular structure, a small-world architecture, a rich-club organisation) are indicators of a genuine degree of self-organisation or simple by-products of lower-order constraints [^11] More explicitly, the practice of employing graph-theoretical estimations for inter-subject analyses without comparing them with a proper benchmark gives ‘[…] a fairly unspecific characterisation of the brain, being fragile to noise […] Another caveat is that […] correlation matrices display small-world properties […] by construction […] This observation highlights the need for well-defined null hypotheses […] but also for controlled recovery of brain functional connectivity going beyond empirical correlation matrices […]’ Varoquaux and Craddock (2013)..
@@
In the present contribution, we focus on the issue of representing sequential data as relational data, by approaching the question how can a set of time series be represented as a traditional graph? from the perspective of traditional hypothesis testing. After comparing the statistical properties of the former against those expected under a suitable benchmark, our suggestion is that of projecting the set of time series onto a signed graph, by linking any two series with a positive (negative) edge in case they share a statistically significant number of concordant (discordant) values. More formally, we adopt an information-theoretic approach that is rooted into the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of their empirical properties on average while randomising everything else Park and Newman (2004); Garlaschelli and Loffredo (2008); Squartini and Garlaschelli (2011).
@@
Our method deals with signed data Gallo et al. (2024a, b, 2025): one may, thus, wonder how relevant is the information provided by signs within the neuroscientific context. Several studies have recently investigated the role played by the negative correlations characterising resting-state fMRI data (rs-fMRI): the authors of Saberi et al. (2021) show that negative correlations are essential for the stability of resting-state brain networks against perturbations or changes; the authors of Saberi et al. (2022) notice that the interplay between positive and negative correlations induces frustrated patterns preventing the brain from reaching a state of minimum energy and enhancing its adaptability to dynamical changes; finally, the authors of Demertzi et al. (2022) show that the default mode network (DMN) works ‘in opposition’ with the various task-positive networks, a supposedly crucial feature to maintain cognitive functions and consciousness.
@@
Let us, however, stress that pre-processing procedures like the Global Signal Regression may artificially induce negative correlations, an evidence further stressing the importance of such a preliminary phase when cleaning neuroscientific data Murphy et al. (2009).
@@ II Setting up the formalism
are the average value and standard deviation of series $i$. Since $w_{it}$ can be either positive or negative, to ease mathematical manipulations let us employ Iverson’s brackets - a notation ensuring all quantities of interest to be non-negative Gallo et al. (2024a) - and define
@@
in an entry-wise fashion - or, more compactly, $\mathbf{B}\equiv\mathbf{B}^{+}-\mathbf{B}^{-}$ and $|\mathbf{B}|\equiv\mathbf{B}^{+}+\mathbf{B}^{-}$. The present analysis solely focuses on binarised time series: beside easing the mathematical tractability of the latter, binarisation has been also shown to lead to better clustering results Thomas Yeo et al. (2011).
Figure 1: Infographics illustrating our analysis pipeline. Pictorial representation of the pipeline we follow in the present contribution, from the registration of brain activity to the identification of communities in statistically validated signed projections: (a) rs-fMRI signals are recorded; (b) pairwise interactions between brain regions are considered; (c) the scalar product of each pair of (standardised and binarised) time series is calculated to quantify their signature; (d) its empirical value is validated against a null model to remove statistical noise; (e) a significantly large positive (negative) signature induces a positive (negative) link between the two involved brain regions; (f) a BIC-based community detection reveals the modules partitioning our statistically validated signed projections.

### v3

@@ I.2 From univariate time series to graphs
Different (classes of) approaches that seek to extract a network from one or more time series while ignoring (the intermediate step represented by) the proximity matrix exist [^9] We redirect the interested reader to the reviews Silva2021; Zou2019, discussing the differences between the existing (classes of) algorithms and comparing their performances..
@@ I.2.1 Visibility-induced mapping
The first algorithm of this kind was proposed in Lacasa2008 and is based on the idea of representing each observation as a vertical bar whose height equals the numerical value of the observation itself: if the tops of any two bars can be directly connected via a ‘visibility line’, the corresponding nodes are, then, connected in the resulting network. The authors of Lacasa2008 have shown that visibility graphs (VGs) inherit several properties from the original time series, e.g. VGs are always connected because each observation (node) is visible by, at least, its temporal (structural) neighbours; hubs are induced by (local) maxima, since the corresponding bars see more nodes than other points; the presence of hubs usually induces a modular structure; a periodic time series induces a network whose degree distribution is characterised by a number of peaks that is smaller than its period $T$, as the degree of node $i$ will be the same of the nodes $i\pm T$, $i\pm 2T$, $i\pm 3T$ and so on; a white noise process induces a (random) network whose degree distribution is exponential while a fractal time series induces a (scale-free) network whose degree distribution is power-law. Visibility algorithms are easy to implement, computationally fast, and parameter-free; besides, they have been employed to diagnose Alzheimer’s disease automatically, achieving an accuracy of $\simeq 98\%$ Ahmadlou2010.
@@
A variant of the aforementioned procedure has been proposed in Luque2009, where the concept of horizontal visibility - as opposed to the one of natural visibility - is defined. A horizontal visibility graph (HVG) is constructed by considering (only) horizontal visibility lines: as a consequence, the HVG is always a subgraph of the VG, hence containing less information, although its construction is faster. HVGs have been employed to analyse seismic signals Telesca2012, the volatility of returns of financial time series Zhang2015, heartbeat rates Xie2019, and for classification tasks Li2018.
@@
A third variant is defined by implementing the concept of directed visibility, whose introduction has been justified by noticing that time has a natural direction: in Lacasa2012, the authors adopted the rule according to which any pair of time steps $t_{i}<t_{j}$ populates a directed horizontal visibility graph (DHVG) with an edge $i\rightarrow j$ and employed DHVGs to quantify time series irreversibility.
@@
A fourth variant is defined by implementing the concept of limited penetrable visibility, proposed in Zhou2012 to reduce the level of noise accompanying data to a larger extent than that achievable by using VGs and HVGs. In this case, any two nodes are connected if the visibility line joining the tops of the two associated bars crosses a number of ‘obstacles’ that is smaller than, or equal to, a fixed value $L$ - naturally, the limited penetrable visibility graph (LPVG) reduces to the VG when $L=0$, i.e. when there are no ‘obstacles’. LPVGs have been employed to analyse EEG signals of people in different conditions (affected by Alzheimer’s disease Wang2016 or under manual acupuncture Pei2014).
@@
A fifth variant is defined by implementing the concept of weighted visibility. The method, proposed in Supriya2016, prescribes to enrich each edge of a DHVG with a weight matching the ‘view angle’ of the observing node, its sign with respect to the horizontal axis indicating if the values of the time series increase (if positive) or decrease (if negative); since such a method allows for negative weights, the absolute value of the angle is often considered. The authors employed the method to identify EEG signals associated with epilepsy.
@@
A last variant is defined by implementing the concept of differential visibility, which is realised by removing the edge set of the HVG from the edge set of the VG Zhu2014.
@@ I.2.2 Transition-induced mapping
The mapping introduced in Campanharo2011 is based on a partition of the distribution of the values assumed by a time series in quantiles, to each of which a symbol is assigned - the number of quantiles is an arbitrary parameter whose choice should depend on the length of the series itself Shirazi2009. The variant of this method introduced in Weng2017 allows ‘memory networks’ to be constructed by assigning the attribute $t$ to the edge directed from $i$ to $j$ if the transition between the corresponding symbols occurs at time step $t$. A second variant is the one introduced in Campanharo2016, where the concept of renewal intervals, i.e. transitions between symbols posed at various temporal distances, is introduced.
The mapping introduced in Gao2009 is named Coarse-Grained Phase Space Graphs and requires the definition of a time window $\omega$ to determine the number of values of the series to be represented as an $\omega$-dimensional vector, whose components are numerically determined by adopting a rule similar to the quantiles-based one.
A different, yet related, mapping is the one introduced in Small2013, named Ordinal Partition Transition Networks. Here, the time series is embedded in a $\omega$-dimensional space where, starting from each point of the series, the current and the following $\omega-1$ numbers, interspaced by a time delay $\tau$, are considered (for instance, setting $\omega=3$ and $\tau=2$, the $i$-th embedding vector will contain the values at $i$, $i+2$, and $i+4$). The embedding vectors are, then, associated with patterns induced by the rank of their values, and each ordinal pattern is represented as a node of the network: nodes $i$ and $j$ are connected if the corresponding patterns occur at subsequent time steps (i.e. if $t_{j}=t_{i}+1$).
A third variant of the same kind of mapping is the one introduced in Mutua2015 and named Visibility Graphlets Networks. This method requires the definition of a time window $\omega$ to determine the number of values to be considered: the time steps from $t$ to $t+\omega$ are, then, mapped into a directed visibility graph (DVG). Such a procedure is repeated for each time step, and the DVGs having the same adjacency matrix are associated with the same symbol, i.e. one of the nodes of the resulting network. Different nodes are linked according to the subsequent occurrences of the related DVGs.
@@ I.2.3 Proximity-induced mapping
The first method of this kind was introduced in Zhang2006 to represent pseudo-periodic time series. First, it segments a time series into cycles (e.g. dictated by a periodic behaviour) without requiring them to have the same length; then, correlations between pairs of cycles are computed - by shifting the shortest one, if needed - and the largest correlation value is compared against a threshold: if found to be larger, an edge between the two cycles, represented as nodes in the resulting network, is established.
An alternative method for employing correlations to implement a proximity-induced mapping is introduced in Yang2008 and prescribes computing the correlations between the (either overlapping or non-overlapping) portions into which a time series can be partitioned: the nodes representing them are then linked if the correlation between the corresponding portions exceeds a certain threshold.
A different class of methods is the one gathering the so-called Recurrence Networks (RNs), which is based on the idea that the points of a time series are recurrent if their distance, in a properly defined embedding space, is ‘short enough’. Similarly to the prescription characterising the Ordinal Partition Transition Networks algorithm, each time step is mapped to a point of an embedding space. Distances are, then, computed by using one of the many norms that can be defined. The way the embedding space is analysed distinguishes RNs in three sub-categories: i) the first one, introduced in Small2009, is composed by the so-called k-Nearest Neighbour Networks and requires to connect each node with the $k$ nodes whose corresponding points in the embedding space are closest to its corresponding point - to be noticed that such a variant has been defined for directed networks; ii) the second one, introduced in Xu2008, is composed by the so-called Adaptive Nearest Neighbour Networks and implements the undirected version of the aforementioned recipe; iii) the third one, introduced in Donner2010, is composed by the so-called $\epsilon$-Recurrence Networks ($\epsilon$-RNs) and prescribes to link the nodes whose corresponding points lie at a distance smaller than $\epsilon$: in symbols, the generic element of an $\epsilon$-RN matrix is defined as
@@ I.3.1 Single-layer mapping
This first kind of mapping collects generalisations of the aforementioned RNs, hereby named Cross Recurrence Networks (CRNs), Inter-System Recurrence Networks (ISRNs), and Joint Recurrence Networks (JRNs) Feldhoff2012; Feldhoff2013; Romano2004.
@@
with $\bm{x}_{t}$ and $\bm{y}_{t^{\prime}}$ indicating the points of the chosen embedding space representing the values of the time series $X_{t}$ and $Y_{t^{\prime}}$ - notice that the two time series do not need to be of the same length, hence inducing a rectangular table Marwan2002; Zbilut1998.
@@
constituted by (matrices representing the) RNs along the diagonal and by (matrices representing the) CRNs off the diagonal Feldhoff2012.
An alternative recipe is represented by the JRNs Romano2004; Feldhoff2013, inducing adjacency matrices like the one reading
@@
A generalisation of the Ordinal Partition Transition Networks for multivariate time series was introduced in Zhang2017. The method requires a binarisation of the univariate time series, letting an increasing trend [^10] Defined as ‘the value of the series at time $t$ is larger than the value of the series at time t1t-1’. being represented with a $1$ and a decreasing trend being represented with a 0: upon doing so, $2^{N}$ combinations, with $N$ being the number of time series, are possible for each time step, $t$; finally, each combination is assigned a node in the resulting network, any two nodes connected by a directed edge representing combinations found at subsequent steps. For deterministic and periodic time series, some of the nodes remain isolated; in contrast, random processes are associated with densely (if not completely) connected graphs.
@@ I.3.2 Multiple-layer mapping
This second kind of mapping, on the other hand, associates a multivariate time series with a multi-layer network. A first example of such a mapping is provided by Lacasa2015: here, a multiplex visibility graph (MVG) is obtained by mapping each univariate time series into a graph by employing a VG; afterwards, the MVG is aggregated into a single-layer weighted graph, each node representing a layer-specific VG and each weight proxying the similarity between the corresponding layers. Such a method has been employed to analyse fMRI data Sannino2017.
The authors of Eroglu2018 follow the same pipeline but replace VGs with RNs, thus employing Multiplex Recurrence Networks (MRNs) to study paleoclimatic time series.
@@ I.4 From multivariate time series to
The most recent approaches emphasise the importance of probabilistic (null, as well as generative) models to benchmark network features and disentangle fundamental, structural properties from statistical artifacts VAROQUAUX2013; vasanull2022. These models play a crucial role in understanding whether empirical features (e.g. a modular structure, a small-world architecture, a rich-club organisation) are indicators of a genuine degree of self-organisation or simple by-products of lower-order constraints [^11] More explicitly, the practice of employing graph-theoretical estimations for inter-subject analyses without comparing them with a proper benchmark gives ‘[…] a fairly unspecific characterisation of the brain, being fragile to noise […] Another caveat is that […] correlation matrices display small-world properties […] by construction […] This observation highlights the need for well-defined null hypotheses […] but also for controlled recovery of brain functional connectivity going beyond empirical correlation matrices […]’ VAROQUAUX2013..
@@
In the present contribution, we focus on the issue of representing sequential data as relational data, approaching the question how can a set of time series be represented as a traditional graph? from the perspective of traditional hypothesis testing. After comparing the statistical properties of the former against those expected under a suitable benchmark, our suggestion is to project the set of time series onto a signed graph by linking any two series with a positive (negative) edge in case they share a statistically significant number of concordant (discordant) values. More formally, we adopt an information-theoretic approach that is rooted in the constrained maximisation of Shannon entropy, a procedure inducing an ensemble of multivariate time series that preserves some of their empirical properties on average while randomising everything else Park2004; Garlaschelli2008; Squartini2011.
@@
As our method deals with signed data gallo2024testing; gallo2024assessing; gallo2025statistically, one may wonder how relevant the information provided by signs is within the neuroscientific framework. Several studies have recently investigated the role played by the negative correlations characterising resting-state fMRI data (rs-fMRI): the authors of Saberi2021 show that negative correlations are essential for the stability of resting-state brain networks against perturbations or changes; the authors of Saberi2022 notice that the interplay between positive and negative correlations induces frustrated patterns preventing the brain from reaching a state of minimum energy and enhancing its adaptability to dynamical changes; finally, the authors of Demertzi2022 show that the default mode network (DMN) works ‘in opposition’ with the various task-based networks, a supposedly crucial feature to maintain cognitive functions and consciousness.
@@
Let us, however, stress that pre-processing procedures, such as Global Signal Regression, may artificially induce negative correlations, further emphasizing the importance of such a preliminary phase when cleaning neuroscientific data Murphy2009.
@@ II Setting up the formalism
are the average value and standard deviation of series $i$. Since $w_{it}$ can be either positive or negative, to ease mathematical manipulations, let us employ Iverson’s brackets - a notation ensuring all quantities of interest are non-negative gallo2024testing - and define
@@
in an entry-wise fashion - or, more compactly, $\mathbf{B}\equiv\mathbf{B}^{+}-\mathbf{B}^{-}$ and $|\mathbf{B}|\equiv\mathbf{B}^{+}+\mathbf{B}^{-}$.
The present analysis solely focuses on binarised time series: besides easing the mathematical tractability of the latter, binarisation has also been shown to lead to better clustering results yeo2011organization. Adopting the terminology introduced for bipartite networks, we will also refer to the rectangular matrix $\mathbf{B}$ as the biadjacency matrix associated with the considered set of time series.
Figure 1: Infographics illustrating the pipeline of our analysis. Pictorial representation of the pipeline we follow in the present contribution, from the registration of brain activity to the identification of communities in statistically validated signed projections: (a) rs-fMRI signals are recorded; (b) pairwise interactions between brain regions are considered; (c) the scalar product of each pair of (standardised and binarised) time series is calculated to quantify their signature; (d) its empirical value is validated against a null model to remove statistical noise; (e) a significantly large positive (negative) signature induces a positive (negative) link between the two involved brain regions; (f) a BIC-based community detection reveals the modules partitioning our statistically validated signed projections. The first brain image was adapted from Canva (additional elements were added by the authors), while the last one was obtained through the Enigma Toolbox Lariviere2021.